# A large data semi-global existence and convergence theorem for vacuum Einstein’s equations

### v3

@@ 1.3. Heuristics on the genericity of the initial data
Corollary 1.1.
Conjecture 1.1 holds in the class of $(3+1)$-dimensional solutions to the vacuum Einstein equations with positive cosmological constant $\Lambda>0$, defined on globally hyperbolic spacetimes $\widetilde{M}\cong M\times\mathbb{R}$ where $M$ is a closed $3$-manifold of negative Yamabe type. The result holds for a class of large initial data prescribed on a Cauchy hypersurface at a Newtonian-like time $T_{0}=a$. In particular, no future naked singularities can form within this setting.
1.4. Comparison with previous studies
The seminal heuristic investigation of the asymptotic dynamics of initially expanding spatially homogeneous cosmological spacetimes governed by Einstein’s field equations with a positive cosmological constant $\Lambda>0$ was undertaken by Wald [48], wherein it was suggested that such solutions generically exhibit de Sitter-like behavior at late times. Subsequently, Ringström [29] established a rigorous framework for analyzing the future development of spatially homogeneous solutions by localizing the global Cauchy problem to a suitable coordinate neighborhood. Exploiting this localization, he proved nonlinear stability and future geodesic completeness for broad classes of spatially homogeneous cosmological models with arbitrary compact spatial topologies.
@@ 1.4. Comparison with previous studies
(10) $\widetilde{g}=\Omega^{2}g,$
supplemented by a hierarchy of conformal curvature variables and auxiliary fields required to ensure hyperbolicity and regularity at the conformal boundary. The initial data are posed not on a standard Cauchy hypersurface but rather at past or future conformal infinity $\mathcal{J}^{\pm}$, where one prescribes a Riemannian metric $\widehat{h}_{ij}$ on the 33–manifold $\mathcal{J}^{\pm}$ (defined only up to an overall conformal scaling) together with a trace–free symmetric 22–tensor $\widehat{d}_{ij}$, subject solely to a divergence constraint with respect to $\widehat{h}$. Crucially, in contrast to the Cauchy formulation, no Hamiltonian (scalar) constraint is imposed at $\mathcal{J}^{\pm}$; the conformal transformation effectively removes the obstruction associated with the scalar constraint at the hypersurface on which $\Omega=0$.
@@
By contrast, the present work addresses a fundamentally distinct regime: it employs a CMC–time slicing on a spatially compact spacetime. The construction accommodates genuinely large data and is intrinsically tied to geometric control of high-order norms within the CMC foliation, rather than to asymptotic smallness near $\mathcal{J}^{\pm}$. In particular, the Hamiltonian and momentum constraints are fully retained and enforced at the initial Cauchy hypersurface, in stark contrast with Friedrich’s boundary data prescription. Furthermore, in the absence of $\Lambda>0$, the conformal extension techniques utilized by Friedrich are, in general, obstructed by the nonexistence of smooth conformal completions of asymptotically simple solutions to the pure vacuum system $\mathrm{Ric}(g)=0$. This obstruction highlights the structural incompatibility between conformal compactification and the non–conformally rescaled large–data interior approach developed in the present analysis.
In contrast, the present study is formulated on a physical compact Cauchy slice at large CMC–time $T_{0}=a$, of negative Yamabe type. We work with the full set of Cauchy data $(g_{0},k_{0})$ satisfying the nonlinear Einstein constraint equations. Furthermore, to the author’s knowledge, Friedrich’s results pertain to an open set of data near de Sitter, typically on manifolds of spherical topology, whereas our construction admits large data consistent with the prescribed CMC hierarchy. In particular, the initial profiles may exhibit strongly concentrated gravitational radiation (reasonably high-frequency), with amplitudes of order $O(e^{a/10})$ supported in regions of spatial extent $O(e^{-a/10})$, while still obeying the required geometric and analytic constraints.
@@
In the present work, we extend this line of inquiry in several significant directions. First, we dispense with the assumption of a homogeneous background solution and instead consider a large class of initial data sets satisfying the Einstein constraint equations on compact manifolds MM of negative Yamabe type, specifically those with $\sigma(M)\leq 0$, as characterized in equation (2). These manifolds rarely admit any locally homogeneous (or isotropic) metrics. Our main result constitutes, to the best of our knowledge, the first rigorous demonstration of the global-in-time well-posedness of solutions to the Einstein equations with $\Lambda>0$, emanating from large, non-perturbative, and reasonably generic initial data on compact manifolds with $\sigma(M)\leq 0$ in dimension $n=3$.
@@ 1.5. Novelty and the main difficulty
In the analysis of nonlinear hyperbolic evolution equations, it is often heuristically observed that small data global existence on $0,\infty)$ may correspond to weighted large data global existence on $T_{0},\infty)$ for sufficiently large $T_{0}$. In the context of Einstein’s equations with positive cosmological constant, one might conjecture that the small data global existence results established in [13, 34] could be extended to yield large data (unweighted) results by considering initial time $T_{0}\gg 1$. This expectation is, however, not valid in the current setting for both geometric and analytic reasons.
To elucidate this failure of transference, consider initial data at time $T_{0}\gg 1$ satisfying
(11) $C^{-1}\xi_{0}\leq g_{0}\leq C\xi_{0},\qquad\sum_{0\leq I\leq 2}\|\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}\leq\epsilon_{0},$
along with high-order control:
(12) $\sum_{0\leq I\leq 3}\|\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}+\sum_{0\leq I\leq 2}\|\nabla^{I}\mathfrak{T}[g_{0}]\|_{L^{2}(M)}\leq\mathcal{I}^{0},$
where $\mathfrak{T}[g]:=\operatorname{Ric}[g]-\frac{1}{n}R(g)g$ is the trace-free Ricci tensor. $\epsilon_{0}$ is sufficiently small and $\mathcal{I}^{0}$ is sufficiently large. Despite the smallness of $\Sigma_{0}$ in low-order norms, the high-order norms of $\Sigma_{0}$ and all norms of $\mathfrak{T}[g_{0}]$ may remain large. As a consequence, the data do not satisfy the hypotheses of small data stability results such as those in [13, 34], and no global control follows.
By contrast, suppose that at $T_{0}=0$, both $\Sigma_{0}$ and $\mathfrak{T}[g_{0}]$ satisfy uniform smallness conditions:
(13) $C^{-1}\xi_{0}\leq g_{0}\leq C\xi_{0},\qquad\sum_{0\leq I\leq 2}\|\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}\leq\epsilon_{0},$
(14) $\sum_{0\leq I\leq 3}\|\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}+\sum_{0\leq I\leq 2}\|\nabla^{I}\mathfrak{T}[g_{0}]\|_{L^{2}(M)}\leq\epsilon_{0},$
for $\epsilon_{0}$ sufficiently small. Then the global existence and nonlinear stability results of [13] are applicable. However, it is a consequence of [52] that these assumptions (along with a lower bound on the eigenvalue of the Lichnerowicz Laplacian) imply that the initial Cauchy hypersurface $(M,g_{0})$ admits an Einstein metric, i.e., $(M,g_{0})$ is hyperbolizable when $n=3$. Thus, the result becomes restricted to small perturbations of spatially homogeneous (e.g., de Sitter) solutions and does not extend to general negative Yamabe manifolds, such as those with graph manifold structure. Therefore, this case can never yield any global control in the previous case when the obstruction tensor $\mathfrak{T}$ and the top order norm of $\Sigma_{0}$ are large. In more physical terms, as stated in section 1.2, the data is large in the sense that the electric part of the Weyl curvature (free gravity) can be arbitrarily large and concentrated on small measure subsets. This is novel and to our knowledge the first result of this kind.
In contrast, the present article develops a framework in which the initial spatial metric $g_{0}$ is arbitrary (subject to geometric bounds), and $\mathfrak{T}[g_{0}]$ is bounded but not required to be small. No Einstein metric structure on the Cauchy hypersurface is assumed. This generality permits global existence and control of large data on non-hyperbolizable topologies, significantly extending the scope of prior stability results.
Secondly, observe that the Hamiltonian constraint associated with the Einstein equations implies that the scalar curvature of the spatial metric $g$ satisfies
(15) $R(g)+\frac{n-1}{n}=|\Sigma|^{2}.$
Under the hypotheses of Theorem 1.1, in particular the bound on $\Sigma$ in $W^{2,2}$, it follows immediately from (15) that
(16) $\left\|R(g)+\frac{n-1}{n}\right\|_{W^{2,2}(M)}\leq e^{-a}\mathcal{I}^{0}.$
For $a\gg 1$ and $\mathcal{I}^{0}$ fixed, or $\mathcal{I}^{0}$ large but $a$ sufficiently large in terms of $\mathcal{I}^{0}$, the right-hand side can be made small. However, this does not imply that $g$ is close to a metric of constant scalar curvature in a strong topology.
Indeed, for $n\geq 3$, the mapping $g\mapsto R(g)$ is a nonlinear second-order differential operator. More precisely, closeness of scalar curvature in $W^{2,2}$ to a negative constant does not control the closeness of the metric $g$ to a reference constant scalar curvature metric $\gamma$ in, say, $W^{4,2}$ or stronger norms. Consequently, even though $R(g)\approx-\frac{n-1}{n}$ in $W^{2,2}$, the metric $g$ itself may be arbitrarily far from any metric of constant scalar curvature $\frac{n-1}{n}$ in any Sobolev norms.
Furthermore, the bounds in Theorem 1.1 only require $g$ to satisfy uniform ellipticity and boundedness of the trace-modified Ricci tensor $\mathfrak{T}[g]:=\operatorname{Ric}[g]-\frac{1}{n}R(g)g$ in $W^{2,2}(M)$, and thus permit large deviations from any background Einstein metric. Therefore, the convergence result established in Theorem 1.1 describes a genuinely large data dynamical regime, in which the initial metric $g_{0}$ may be far in $W^{4,2}(M)$ (and higher topologies) from any constant scalar curvature or Einstein geometry (even if it exists on $M$). Nevertheless, the solution converges globally in time to a limiting metric of constant scalar curvature.
This phenomenon—global convergence to a constant scalar curvature geometry from arbitrary (non-small) initial data—is not captured by previous small perturbation results around Einstein metrics (cf. [13]). To the best of the author’s knowledge, this work constitutes the first rigorous treatment of the large data convergence problem for the Einstein equations with positive cosmological constant on general compact spatial manifolds in dimensions $n\geq 3$.
We now elaborate on a key structural aspect of our formulation. Rather than employing the coupled hyperbolic-elliptic system for $(g,\Sigma)$ in constant mean curvature (CMC) and spatial harmonic (SH) gauge, we adopt the manifestly hyperbolic evolution system for $(\Sigma,\mathfrak{T})$, where $\mathfrak{T}_{ij}:=\operatorname{Ric}_{ij}-\frac{1}{n}R(g)g_{ij}$, coupled to an elliptic lapse equation (76), and impose the CMC-transported spatial coordinate gauge (i.e., vanishing shift vector field). This formulation is motivated by the classical work of Choquet-Bruhat and York [1], who studied a hyperbolic system for $(\Sigma,\operatorname{Ric})$. In this gauge, the evolution equation for $g$ (cf. (64)) is treated as a transport equation along the future-directed unit normal vector field $\partial_{T}$, whose flow lines form the characteristics. In contrast, under spatial harmonic gauge, the metric evolution is governed by a quasilinear wave equation with null cone characteristics. More significantly, the SH gauge introduces serious analytical complications in the large data regime, which we now make precise.
Preservation in time of the SH gauge condition $V^{i}=0$, where $V^{i}:=g^{jk}(\Gamma[g]^{i}_{jk}-\Gamma[\xi]^{i}_{jk})$ is the tension field associated to a fixed background metric $\xi$, yields the following elliptic equation for the shift vector field $X$:
(17) $\mathcal{P}X^{i}:=-\Delta_{g}X^{i}-\operatorname{Ric}^{i}_{j}[g]X^{j}+2\nabla^{j}X^{k}(\Gamma[g]^{i}_{jk}-\Gamma[\xi]^{i}_{jk})$ $=\mathcal{F}^{i}[g,\Sigma,N],$
where $\mathcal{F}^{i}[g,\Sigma,N]$ denotes a source term depending on the lapse $N$, the second fundamental form $\Sigma$, and $g$ relative to $\xi$ (cf. [13, 34]). In previous small data studies [13, 34], $\xi$ is chosen to be a fixed hyperbolic metric and $g$ is assumed close to $\xi$ in $W^{s,2}$ with $s>\frac{5}{2}$, so that $\mathcal{P}$ defines an isomorphism $W^{s+1,2}\to W^{s-1,2}$.
Let us sketch the standard argument in the small data regime. The operator $\mathcal{P}$ is elliptic and may be decomposed as
(18) $\mathcal{P}X^{i}+\mathscr{L}_{X}V^{i}$ $=\Delta^{\xi}_{g}X^{i}-g^{k\ell}R[\xi]^{i}{}_{kj\ell}X^{j},$
where $\mathscr{L}_{X}V$ denotes the Lie derivative of the tension field, and $\Delta^{\xi}_{g}$ is the rough Laplacian defined using $\xi$ and acting on vector fields (cf.  [3]). The derivation follows from differentiating the pullback identity
$(\psi_{s}^{*}V)^{i}=(\psi_{s}^{*}g^{-1})^{jk}(\Gamma[\psi_{s}^{*}g]^{i}_{jk}-\Gamma[\psi_{s}^{*}\xi]^{i}_{jk}),$
where $\psi_{s}$ is the flow of $X$. In particular, under the smallness assumption $\|g-\xi\|_{W^{s,2}}\ll 1$, the operator $\mathcal{P}_{g,\xi}:=\Delta^{\xi}_{g}-g^{k\ell}R[\xi]^{i}{}_{kj\ell}$ becomes coercive. Indeed, integration by parts yields:
$0=\int_{M}\xi_{ij}Z^{i}\mathcal{P}_{g,\xi}Z^{j}\,d\mu_{\xi}$ $=\int_{M}\left[\xi_{ij}\xi^{k\ell}\nabla_{k}Z^{j}\nabla_{\ell}Z^{i}-\xi_{im}\xi^{k\ell}Z^{m}R[\xi]^{i}{}_{kj\ell}Z^{j}+\text{Err}_{g,\xi}\right]d\mu_{\xi},$
where the error terms $\text{Err}_{g,\xi}$ are perturbative if $\|g-\xi\|_{W^{s,2}}$ is small. Since $R[\xi]^{i}{}_{j}$ is negative-definite (in the case $\xi$ is hyperbolic), the integrand is strictly positive unless $Z\equiv 0$, hence $\ker\mathcal{P}_{g,\xi}=\{0\}$ and the operator is an isomorphism (note it is self-adjoint and elliptic).
However, this strategy fails in the large data regime. In particular, the general Cauchy hypersurfaces considered here are compact, but may carry no hyperbolic metric at all, for example when $\sigma(M)\leq 0$. Moreover, in the absence of smallness assumptions on $\|g-\xi\|_{W^{s,2}}$, the leading-order perturbative structure in the elliptic operator $\mathcal{P}$ degenerates, and both the Ricci and first-order connection terms in (17) are no longer controlled. This introduces Gribov-type ambiguities in solving for $X$, akin to gauge degeneracy in non-Abelian Yang-Mills theory in Coulomb gauge [18]. Indeed, the SH gauge is analogous to the Coulomb gauge in Yang–Mills theory, where existence and uniqueness of global solutions is obstructed by the presence of Gribov horizons.
In contrast, the CMC-transported spatial coordinate gauge circumvents this obstruction entirely: the shift vector field is identically zero, and the evolution system is strictly hyperbolic modulo a single elliptic lapse equation. This choice parallels the temporal gauge in Yang–Mills theory, which avoids the Gribov ambiguity but introduces technical subtleties in constraint propagation. Our analysis reveals that in the gravitational setting, the CMC-transported spatial gauge provides a natural framework for addressing large data evolution problems without encountering elliptic degeneracy in gauge fixing.
A key technical ingredient in the analysis is the construction of a hierarchy of energy norms for the transverse-traceless second fundamental form $\Sigma$. These norms are designed to exploit the decay induced by the cosmological expansion. Specifically, although the top-order energy norm of $\Sigma$ does not decay (as reflected in Theorem 1.1), the lower-order norms exhibit exponential decay, which is systematically leveraged in a bootstrap framework to gain successive improvements at each level of regularity. This hierarchy is essential for closing the top-order estimates.
The structure of the Einstein equations plays a crucial role in controlling nonlinear interactions. In particular, the terms that lack decay at top order always appear with an exponentially decaying prefactor $e^{-T}$. While this is not a manifestation of a null structure, it indicates a hidden weak null-type behavior that remains to be fully understood. Furthermore, although the normalized lapse perturbation $\frac{N}{n}-1$ does not decay at top order, all instances where this quantity appears differentiated at the top order are accompanied by a decaying factor $e^{-T}$, while in undamped terms it arises only through lower order derivatives, which ensures integrable decay. We elaborate on this more in section 1.6
All energy and elliptic norms are defined relative to the evolving spatial metric $g$, and hence a priori estimates for $g$ are indispensable. To control the metric, we integrate the evolution equation
along the CMC-transported foliation. As $\Sigma$ is controlled in $W^{I,2}$, this yields control over $g$ in $W^{I,2}$, but not in $W^{I+1,2}$. Consequently, standard elliptic regularity theory, which would require $g\in W^{I+1,2}$ to control $\frac{N}{n}-1$ in $W^{I+2,2}$, does not directly apply.
This leads to an apparent regularity obstruction in estimating the normalized lapse from its governing elliptic equation:
(20) $\Delta_{g}\left(\frac{N}{n}-1\right)+\left(|\Sigma|^{2}+\frac{1}{n}\right)\left(\frac{N}{n}-1\right)=-|\Sigma|^{2}.$
If Sobolev norms were defined relative to a fixed background metric $\xi_{0}$, as in [3], the operator on the left-hand side would define an isomorphism between appropriate Sobolev spaces, yielding the estimate
(21) $\left\|\frac{N}{n}-1\right\|_{W^{I+2,2}(M)}\lesssim\|\Sigma\|^{2}_{W^{I,2}(M)},\qquad I>\tfrac{3}{2}.$
However, in the present setting, we do not have access to the necessary $W^{I+1,2}$ regularity of $g$, either via transport or elliptic estimates, due to the absence of spatial harmonic gauge.
To circumvent this, we define Sobolev norms with respect to the dynamical metric $g$ and commute the elliptic equation (20) with covariant derivatives associated to the Levi-Civita connection of $g$. Upon commuting, curvature terms arise via standard identities, and elliptic regularity for (20) then requires control over the $W^{I-1,2}$ norm of the Riemann curvature of $g$ to estimate $\frac{N}{n}-1$ in $W^{I+2,2}$. One subtlety here is that we need to control the metric components separately through transport equation.
This is resolved through the algebraic decomposition of the Riemann tensor in terms of the electric Weyl tensor $\mathfrak{T}$ and scalar curvature $R[g]$. The latter is in turn expressed via the Hamiltonian constraint:
(22) $R[g]+\frac{n-1}{n}=|\Sigma|^{2}.$
Hence, the Riemann tensor is controlled in $W^{I-1,2}$ in terms of the $W^{I-1,2}$ norms of $\Sigma$ and $\mathfrak{T}$. This suffices to control all curvature-dependent commutator terms in the elliptic theory.
Finally, since the wave equations require $W^{I,2}$ control of $\Sigma$ and $W^{I-1,2}$ control of $\mathfrak{T}$, the above argument provides the needed $W^{I-1,2}$ regularity of the Riemann tensor. Consequently, elliptic regularity theory yields (21), thereby closing the necessary estimates for the lapse.
A further distinguishing feature of this work is the generality of the class of spatial topologies to which our analysis applies. Specifically, we establish our semi-global well-posedness result for every smooth, closed, connected, oriented 3-manifold that admits a metric of negative Yamabe type, i.e., one for which no conformal class contains a metric of non-negative scalar curvature. Equivalently, this includes all manifolds falling into Classes 1–3 in the standard topological classification of closed 3-manifolds. To our knowledge, no prior analysis of the full Einstein equations—particularly at the level of large data energy estimates—has been carried out for such a topologically general class of Cauchy hypersurfaces. Lastly, a major novel feature of this study is sketched out in section 1.3.

### v4

@@ 1.4. Heuristics for the construction of the initial data
Corollary $1.$
Conjecture 1.1 holds in the class of $(3+1)$-dimensional solutions to the vacuum Einstein equations with positive cosmological constant $\Lambda>0$, defined on globally hyperbolic spacetimes $\widetilde{M}\cong M\times\mathbb{R}$ where $M$ is a closed 33-manifold of negative Yamabe type. The result holds for a class of large initial data prescribed on a Cauchy hypersurface at a Newtonian-like time $T_{0}=a$. In particular, no future naked singularities can form within this setting.
1.5. Comparison with previous work
The mathematical analysis of Einstein’s equations with positive cosmological constant has a substantial history, but the regimes treated in the literature differ markedly from the one considered here.
At the level of spatially homogeneous cosmologies, Wald’s cosmic no-hair theorem shows that initially expanding homogeneous solutions with $\Lambda>0$ are driven toward de Sitter behavior at late times [48]. This result is foundational, but it concerns homogeneous dynamics and does not address the large-data Cauchy problem on a general compact spatial manifold. In a different direction, Ringström established future nonlinear stability results for expanding solutions in related Einstein–matter models; in particular, the Einstein equations with positive cosmological constant arise there as a distinguished special case of the more general framework [29]. These works provide essential background for the subject, but they do not furnish the type of compact, large-data, CMC-based theorem proved here.
@@ 1.5. Comparison with previous work
The present theorem operates in a different regime. First, no Einstein background is fixed in advance. The spatial manifold $M$ is only assumed to be a closed 33-manifold with
$\sigma(M)\leq 0,$
and the theorem does not require that $M$ admit a negative Einstein metric. Second, the result is formulated for physical Cauchy data $(g_{0},k_{0})$ posed on a compact slice at time $T_{0}=a$, and the constraint equations are solved explicitly for the class of data under consideration. Third, the theorem is genuinely non-perturbative from the viewpoint of the evolution variables: the size parameter $\mathcal{I}^{0}$ is arbitrary, and the only smallness requirement is the relative condition
$\mathcal{I}^{0}e^{-a/10}<1.$
In particular, the renormalized trace-free Ricci tensor $\mathfrak{T}[g_{0}]$ is not assumed to be small in $H^{2}$, and the unweighted top-order norm of the trace-free second fundamental form $\Sigma_{0}$ is not assumed to be small in $H^{3}$. This places the theorem outside the perturbative framework of [13, 34].
@@
To the best of our knowledge, this is the first future-global well-posedness and convergence theorem for the $3+1$-dimensional vacuum Einstein equations with $\Lambda>0$ on compact CMC slices, based on an explicitly constructed class of physical Cauchy data, that is not perturbative around a prescribed Einstein metric. In particular, it yields global future control and smooth convergence for a nontrivial large-data class on compact 33-manifolds with $\sigma(M)\leq 0$.